OpenStudy (anonymous):

Fun Land theme park has a total of 110 males and females who are employed at the park. If the number of males is 10 more than three times the number of females, how many males are employed at the park?

OpenStudy (anonymous):

set up a system let m = number of males employed at Fun Land let f = number of females employed at Fun Land so your equations are m+f = 110 m= 3f + 10 this is super easy. just substitute in m with (3f+10) so... m+f = 110 (3f+10) + f = 110 4f +10 = 110 4f = 100 f = 25 so there are only 25 females who work at Fun Land now, to find how many males there are, you just need to sub the female value back into the second equation. m = 3(25) + 10 m = 75 + 10 m = 85 so there are 85 males who work at the Fun Land. (if you check your answers, 85 + 25 = 110, so this is correct). do you understand?